If 3 hens lay 4 eggs in 5 days, how many days will it take a dozen hens to lay 8 dozen eggs/
Answer by stanbon(48516)

[(4 eggs/3 hens)/5 days] = [(96 eggs/12 hens)/x days]
[4/15] = [96/12x]
[4/15] = [8/x]
4x = 120
x = 30 days
=====================
Cheers,
Stan H.